# Information conservation in de Sitter tunneling

### v1

@@ 3.2 Tunneling out of the black hole outer event horizon
For the same reason as in Section 2, the WKB approximation can be applied. The emission rate of particles has the same relationship as (61):
@@
In (147), (148) and (150), conservation of energy, conservation of angular momentum, conservation of electric charge and conservation of magnetic charge are manifested. Substitute (147), (148), (149), (150) into (146), we can obtain:
@@
Where $\Delta_{r}=-\frac{\Lambda}{3}(r-r_{non})(r-r_{-})(r-r_{+})(r-r_{c})$. Substitute (152) into (151), we can obtain:
$\mathrm{Im}S=$ $-\mathrm{Im}\{\int_{r_{i}}^{r_{f}}[\int_{(0,0,0)}^{(\omega,q_{h},j)}\frac{6\Xi^{\prime}}{\Lambda\Sigma^{\prime}}\frac{1}{(r-r_{non}^{\prime})(r-r_{-}^{\prime})(r-r_{+}^{\prime})(r-r_{c}^{\prime})}\times$
$\sqrt{\frac{(\Delta_{r}^{\prime}-\Sigma^{\prime})\Sigma^{\prime}(h^{\prime 2}\Delta_{r}^{\prime}-\xi^{\prime 2}\Delta_{\theta}^{\prime}\sin^{2}\theta)}{\Delta_{\theta}^{\prime}\sin^{2}\theta}}\left(\mathrm{d}\omega^{\prime}-\frac{(Q_{h}-q_{h}^{\prime})r}{r^{2}+a^{\prime 2}}\mathrm{d}q^{\prime}_{h}-\Omega^{\prime}\mathrm{d}j^{\prime}\right)]\mathrm{d}r\}.$ (153)
@@
By substituting (155) into (154), we get our final result:
@@
(157) indicates that our results conform to the unitary principle and support the conservation of information.
@@ 3.3 Tunneling into the cosmic horizon
(160), (161) and (163) clearly represent conservation of energy, conservation of angular momentum, conservation of charge and conservation of magnetic charge. Substitute (158), (160), (161), (162) and (163) into (159), we can obtain:
$\mathrm{Im}S=$ $-\mathrm{Im}\{\int_{r_{i}^{\prime}}^{r_{f}^{\prime}}[\int_{(0,0,0)}^{(\omega,q_{h},j)}\frac{6\Xi^{\prime\prime}}{\Lambda\Sigma^{\prime\prime}}\frac{1}{(r-r_{non}^{\prime\prime})(r-r_{-}^{\prime\prime})(r-r_{+}^{\prime\prime})(r-r_{c}^{\prime\prime})}\times$
$\sqrt{\frac{(\Delta_{r}^{\prime\prime}-\Sigma^{\prime\prime})\Sigma^{\prime\prime}(h^{\prime\prime 2}\Delta_{r}^{\prime\prime}-\xi^{\prime\prime 2}\Delta_{\theta}^{\prime\prime}\sin^{2}\theta)}{\Delta_{\theta}^{\prime\prime}\sin^{2}\theta}}\left(\mathrm{d}\omega^{\prime}-\frac{(Q_{h}+q_{h}^{\prime})r}{r^{2}+a^{\prime\prime 2}}\mathrm{d}q^{\prime}_{h}-\Omega^{\prime}\mathrm{d}j^{\prime}\right)]\mathrm{d}r\}.$ (164)
@@
By substituting (166) into (165), we get our final result:
@@
Clearly, (168) indicates that our results conform to the unitary principle and support the conservation of information.
@@ 3.4 Discussion
Comparing (170) and (171) with (154) and (165) respectively, the temperature $T^{\prime}$ at the event horizon surface of the black hole and the temperature $T^{\prime\prime}$ at the cosmic horizon surface are obtained:
@@ 4 Bardeen de Sitter black hole
It is evident from (175) that the electromagnetic field in Bardeen de Sitter spacetime is nonlinear. Therefore, it is impossible to prove the conservation of information by the concrete form of the horizon $r_{h}(M,Q_{h})$. In this section, we use the first law of thermodynamics of black holes to prove the conservation of information. According to the basic knowledge of black hole thermodynamics, the Hawking temperature for spherically symmetric black holes has the following expression:
@@ 4.1 Painlevé coordinate and time-like geodesic line equation
The positive sign represents particles exiting from the black hole event horizon, and the negative sign represents particles entering from the cosmic horizon. Considering the effect of self-gravitation, (179) and (180) should be replaced as follows, $M\to M\mp\omega$, $Q_{h}\to Q_{h}\mp q_{h}$. The minus sign and plus sign represent outgoing particles and incident particles, respectively.
@@ 4.2 Tunneling out of the black hole outer event horizon
$\dot{r}=\frac{\mathrm{d}H}{\mathrm{d}P_{r}}\bigg{|}_{(r;\tilde{A}_{0},P_{\tilde{A}_{0}})}=\frac{\mathrm{d}(M-\omega^{\prime})}{\mathrm{d}P_{r}}=-\frac{\mathrm{d}\omega^{\prime}}{\mathrm{d}P_{r}}.$ (185)
@@
In (185) and (186), energy conservation, charge conservation, and magnetic charge conservation are all reflected. By substituting (185) and (186) into (184), we can obtain:
$\mathrm{Im}S=-\mathrm{Im}\left\{\int_{r_{i}}^{r_{f}}\left[\int_{(0,0)}^{(\omega,q_{h})}\frac{1}{\dot{r}}\left(\mathrm{d}\omega^{\prime}-\frac{Q_{h}-q_{h}^{\prime}}{r}\mathrm{d}q_{h}^{\prime}\right)\right]\right\}.$ (187)
@@
Substituting (188) into (187) and considering the influence of self gravity, we can make the following substitution, $M\to M-\omega$, $Q_{h}\to q_{h}^{\prime}$. We can obtain:
$\mathrm{Im}S=-\mathrm{Im}\left\{\int_{r_{i}}^{r_{f}}\left[\int_{(0,0)}^{(\omega,q_{h})}\frac{2r}{f(r)}\sqrt{\frac{2(M-\omega^{\prime})}{[r^{2}+(Q_{h}-q_{h}^{\prime})^{2}]^{3/2}}+\frac{\Lambda}{3}}\left(\mathrm{d}\omega^{\prime}-\frac{Q_{h}-q_{h}^{\prime}}{r}\mathrm{d}q_{h}^{\prime}\right)\right]\right\}.$ (189)
@@
$\mathrm{Im}S=$ $-\int_{(0,0)}^{(\omega,q_{h})}\frac{2\pi}{f^{\prime}(r_{+}^{\prime})}r_{+}^{\prime}\sqrt{\frac{1-f^{(}r_{+}^{\prime})}{r_{+}^{\prime 2}}}\left(\mathrm{d}\omega^{\prime}-\frac{Q_{h}-q_{h}^{\prime}}{r_{+}^{\prime}}\mathrm{d}q^{\prime}_{h}\right)$
$=$ $-\frac{1}{2}\int_{(M,E_{Q_{h}})}^{(M-\omega,E_{Q_{h}-q_{h}})}\frac{1}{T^{\prime}}\left(\mathrm{d}M^{\prime}-\frac{Q_{h}^{\prime}}{r_{+}^{\prime}}\mathrm{d}Q^{\prime}_{h}\right)=-\frac{1}{2}\int_{r_{i}}^{r_{f}}\mathrm{d}S_{BH}^{\prime}=-\frac{1}{2}\Delta S_{BH}.$ (190)
@@ 4.3 Tunneling into the cosmic horizon
In (194) and (195), energy conservation, charge conservation, and magnetic charge conservation are all reflected. By substituting (192), (194) and (195) into (193), we can obtain:
$\mathrm{Im}S=-\mathrm{Im}\left\{\int_{r_{i}}^{r_{f}}\left[\int_{(0,0)}^{(\omega,q_{h})}\frac{2r}{f(r)}\sqrt{\frac{2(M+\omega^{\prime})}{[r^{2}+(Q_{h}+q_{h}^{\prime})^{2}]^{3/2}}+\frac{\Lambda}{3}}\left(\mathrm{d}\omega^{\prime}-\frac{Q_{h}+q_{h}^{\prime}}{r}\mathrm{d}q_{h}^{\prime}\right)\right]\right\}.$ (196)
@@
$\mathrm{Im}S=$ $-\int_{(0,0)}^{(\omega,q_{h})}\frac{2\pi}{f^{\prime}(r_{c}^{\prime\prime})}\left(\mathrm{d}\omega^{\prime}-\frac{Q_{h}+q_{h}^{\prime}}{r_{c}^{\prime\prime}}\mathrm{d}q_{h}^{\prime}\right)$
$=$ $-\frac{1}{2}\int_{(M,E_{Q_{h}})}^{(M+\omega,E_{Q_{h}+q_{h}})}\frac{1}{T^{\prime\prime}}\left(\mathrm{d}M^{\prime}-\frac{Q_{h}^{\prime}}{r_{c}^{\prime\prime}}\mathrm{d}Q_{h}^{\prime}\right)=-\frac{1}{2}\int_{r_{i}^{\prime}}^{r_{f}^{\prime}}\mathrm{d}S_{CH}^{\prime}=-\frac{1}{2}\Delta S_{CH}.$ (197)
Where $\Delta S_{CH}=S_{CH}(M+\omega,Q_{e}+q_{e},Q_{g}+q_{g})-S_{CH}(M,Q_{e},Q_{g})$ is the change in entropy before and after particle enters the cosmic horizon. So the emission rate of particles can be expressed by the following equation:
@@
(198) indicates that our results conform to the unitary principle and support the conservation of information.

### v2

@@ 3.2 Tunneling out of the black hole outer event horizon
For the same reason as in Section 2, the WKB approximation can be applied. The emission rate of particles has the same relationship as Eq. (55):
@@
In Eq. (135), Eq. (136) and Eq. (138), conservation of energy, conservation of angular momentum, conservation of electric charge and conservation of magnetic charge are manifested. Substitute Eq. (135), Eq. (136), Eq. (137), Eq. (138) into Eq. (134), we can obtain:
@@
Where $\Delta_{r}=-\frac{\Lambda}{3}(r-r_{non})(r-r_{-})(r-r_{+})(r-r_{c})$. Substitute Eq. (140) into Eq. (139), we can obtain:
$\mathrm{Im}S=$ $-\mathrm{Im}\left\{\int_{r_{i}}^{r_{f}}\left[\int_{(0,0,0)}^{(\omega,q_{h},j)}\frac{6\Xi^{\prime}}{\Lambda\Sigma^{\prime}}\frac{1}{(r-r_{non}^{\prime})(r-r_{-}^{\prime})(r-r_{+}^{\prime})(r-r_{c}^{\prime})}\times\right.\right.$
$\left.\left.\sqrt{\frac{(\Delta_{r}^{\prime}-\Sigma^{\prime})\Sigma^{\prime}(h^{\prime 2}\Delta_{r}^{\prime}-\xi^{\prime 2}\Delta_{\theta}^{\prime}\sin^{2}\theta)}{\Delta_{\theta}^{\prime}\sin^{2}\theta}}\left(\mathrm{d}\omega^{\prime}-\frac{(Q_{h}-q_{h}^{\prime})r}{r^{2}+a^{\prime 2}}\mathrm{d}q^{\prime}_{h}-\Omega^{\prime}\mathrm{d}j^{\prime}\right)\right]\mathrm{d}r\right\}.$ (141)
@@
By substituting Eq. (143) into Eq. (142), we get our final result:
@@
Eq. (145) indicates that our results conform to the unitary principle and support the conservation of information.
@@ 3.3 Tunneling into the cosmic horizon
Eq. (148), Eq. (149) and Eq. (151) clearly represent conservation of energy, conservation of angular momentum, conservation of charge and conservation of magnetic charge. Substitute Eq. (146), Eq. (148), Eq. (149), Eq. (150) and Eq. (151) into Eq. (147), we can obtain:
$\mathrm{Im}S=$ $-\mathrm{Im}\left\{\int_{r_{i}^{\prime}}^{r_{f}^{\prime}}\left[\int_{(0,0,0)}^{(\omega,q_{h},j)}\frac{6\Xi^{\prime\prime}}{\Lambda\Sigma^{\prime\prime}}\frac{1}{(r-r_{non}^{\prime\prime})(r-r_{-}^{\prime\prime})(r-r_{+}^{\prime\prime})(r-r_{c}^{\prime\prime})}\times\right.\right.$
$\left.\left.\sqrt{\frac{(\Delta_{r}^{\prime\prime}-\Sigma^{\prime\prime})\Sigma^{\prime\prime}(h^{\prime\prime 2}\Delta_{r}^{\prime\prime}-\xi^{\prime\prime 2}\Delta_{\theta}^{\prime\prime}\sin^{2}\theta)}{\Delta_{\theta}^{\prime\prime}\sin^{2}\theta}}\left(\mathrm{d}\omega^{\prime}-\frac{(Q_{h}+q_{h}^{\prime})r}{r^{2}+a^{\prime\prime 2}}\mathrm{d}q^{\prime}_{h}-\Omega^{\prime}\mathrm{d}j^{\prime}\right)\right]\mathrm{d}r\right\}.$ (152)
@@
By substituting Eq. (154) into Eq. (153), we get our final result:
@@
Clearly, Eq. (156) indicates that our results conform to the unitary principle and support the conservation of information.
@@ 3.4 Discussion
Comparing Eq. (158) and Eq. (159) with Eq. (142) and Eq. (153) respectively, the temperature $T^{\prime}$ at the event horizon surface of the black hole and the temperature $T^{\prime\prime}$ at the cosmic horizon surface are obtained:
@@ 4 Bardeen de Sitter black hole
It is evident from Eq. (163) that the electromagnetic field in Bardeen de Sitter spacetime is nonlinear. Therefore, it is impossible to prove the conservation of information by the concrete form of the horizon $r_{h}(M,Q_{h})$. In this section, we use the first law of thermodynamics of black holes to prove the conservation of information. According to the basic knowledge of black hole thermodynamics, the Hawking temperature for spherically symmetric black holes has the following expression:
@@ 4.1 Painlevé coordinate and time-like geodesic line equation
The positive sign represents particles exiting from the black hole event horizon, and the negative sign represents particles entering from the cosmic horizon. Considering the effect of self-gravitation, Eq. (167) and Eq. (168) should be replaced as follows, $M\to M\mp\omega$, $Q_{h}\to Q_{h}\mp q_{h}$. The minus sign and plus sign represent outgoing particles and incident particles, respectively.
@@ 4.2 Tunneling out of the black hole outer event horizon
$\dot{r}=\frac{\mathrm{d}H}{\mathrm{d}P_{r}}\bigg{|}_{(r;\tilde{A}_{0},P_{\tilde{A}_{0}})}=\mathcal{W}\frac{\mathrm{d}(M-\omega^{\prime})}{\mathrm{d}P_{r}}=-\mathcal{W}\frac{\mathrm{d}\omega^{\prime}}{\mathrm{d}P_{r}}.$ (173)
@@
Where $\mathcal{W}=1+\int_{r_{h}}^{+\infty}4\pi r^{2}\frac{\partial T^{0}_{0}}{\partial M}\mathrm{d}r$, for regular black holes, $M$ no longer represents the internal energy, but differs from the internal energy by a factor $\mathcal{W}$. In Eq. (173) and Eq. (174), energy conservation, charge conservation, and magnetic charge conservation are all reflected. By substituting Eq. (173) and Eq. (174) into Eq. (172), we can obtain:
$\mathrm{Im}S=-\mathrm{Im}\left\{\int_{r_{i}}^{r_{f}}\left[\int_{(0,0)}^{(\omega,q_{h})}\frac{1}{\dot{r}}\left(\mathcal{W}\mathrm{d}\omega^{\prime}-\frac{Q_{h}-q_{h}^{\prime}}{r}\mathrm{d}q_{h}^{\prime}\right)\right]\right\}.$ (175)
@@
Substituting Eq. (176) into Eq. (175) and considering the influence of self gravity, we can make the following substitution, $M\to M-\omega$, $Q_{h}\to q_{h}^{\prime}$. We can obtain:
$\mathrm{Im}S=-\mathrm{Im}\left\{\int_{r_{i}}^{r_{f}}\left[\int_{(0,0)}^{(\omega,q_{h})}\frac{2r}{f(r)}\sqrt{\frac{2(M-\omega^{\prime})}{[r^{2}+(Q_{h}-q_{h}^{\prime})^{2}]^{3/2}}+\frac{\Lambda}{3}}\left(\mathcal{W}\mathrm{d}\omega^{\prime}-\frac{Q_{h}-q_{h}^{\prime}}{r}\mathrm{d}q_{h}^{\prime}\right)\right]\right\}.$ (177)
@@
$\mathrm{Im}S=$ $-\int_{(0,0)}^{(\omega,q_{h})}\frac{2\pi}{f^{\prime}(r_{+}^{\prime})}r_{+}^{\prime}\sqrt{\frac{1-f^{(}r_{+}^{\prime})}{r_{+}^{\prime 2}}}\left(\mathcal{W}\mathrm{d}\omega^{\prime}-\frac{Q_{h}-q_{h}^{\prime}}{r_{+}^{\prime}}\mathrm{d}q^{\prime}_{h}\right)$
$=$ $-\frac{1}{2}\int_{(M,E_{Q_{h}})}^{(M-\omega,E_{Q_{h}-q_{h}})}\frac{1}{T^{\prime}}\left(\mathcal{W}\mathrm{d}M^{\prime}-\frac{Q_{h}^{\prime}}{r_{+}^{\prime}}\mathrm{d}Q^{\prime}_{h}\right)=-\frac{1}{2}\int_{r_{i}}^{r_{f}}\mathrm{d}S_{BH}^{\prime}=-\frac{1}{2}\Delta S_{BH}.$ (178)
For regular black holes, the first law of black hole thermodynamics also requires correction with the factor $\mathcal{W}$, so the factor $\mathcal{W}$ in Eq. (178) disappears. Where $\Delta S_{BH}=S_{BH}(M-\omega,Q_{e}-q_{e},Q_{g}-q_{g})-S_{BH}(M,Q_{e},Q_{g})$ is the change in entropy before and after the black hole radiates particles. So the emission rate of particles can be expressed by the following equation:
@@ 4.3 Tunneling into the cosmic horizon
$\dot{r}=\frac{\mathrm{d}H}{\mathrm{d}P_{r}}\bigg{|}_{(r;\tilde{A}_{0},P_{\tilde{A}_{0}})}=\mathcal{W}\frac{\mathrm{d}(M+\omega^{\prime})}{\mathrm{d}P_{r}}=\mathcal{W}\frac{\mathrm{d}\omega^{\prime}}{\mathrm{d}P_{r}}.$ (182)
@@
For the same reason, factor $\mathcal{W}$ will also appear here. In Eq. (182) and Eq. (183), energy conservation, charge conservation, and magnetic charge conservation are all reflected. By substituting Eq. (180), Eq. (182) and Eq. (183) into Eq. (181), we can obtain:
$\mathrm{Im}S=-\mathrm{Im}\left\{\int_{r_{i}}^{r_{f}}\left[\int_{(0,0)}^{(\omega,q_{h})}\frac{2r}{f(r)}\sqrt{\frac{2(M+\omega^{\prime})}{[r^{2}+(Q_{h}+q_{h}^{\prime})^{2}]^{3/2}}+\frac{\Lambda}{3}}\left(\mathcal{W}\mathrm{d}\omega^{\prime}-\frac{Q_{h}+q_{h}^{\prime}}{r}\mathrm{d}q_{h}^{\prime}\right)\right]\right\}.$ (184)
@@
$\mathrm{Im}S=$ $-\int_{(0,0)}^{(\omega,q_{h})}\frac{2\pi}{f^{\prime}(r_{c}^{\prime\prime})}\left(\mathcal{W}\mathrm{d}\omega^{\prime}-\frac{Q_{h}+q_{h}^{\prime}}{r_{c}^{\prime\prime}}\mathrm{d}q_{h}^{\prime}\right)$
$=$ $-\frac{1}{2}\int_{(M,E_{Q_{h}})}^{(M+\omega,E_{Q_{h}+q_{h}})}\frac{1}{T^{\prime\prime}}\left(\mathcal{W}\mathrm{d}M^{\prime}-\frac{Q_{h}^{\prime}}{r_{c}^{\prime\prime}}\mathrm{d}Q_{h}^{\prime}\right)=-\frac{1}{2}\int_{r_{i}^{\prime}}^{r_{f}^{\prime}}\mathrm{d}S_{CH}^{\prime}=-\frac{1}{2}\Delta S_{CH}.$ (185)
For regular black holes, the first law of black hole thermodynamics also requires correction with the factor $\mathcal{W}$, so the factor $\mathcal{W}$ in Eq. (185) disappears. Where $\Delta S_{CH}=S_{CH}(M+\omega,Q_{e}+q_{e},Q_{g}+q_{g})-S_{CH}(M,Q_{e},Q_{g})$ is the change in entropy before and after particle enters the cosmic horizon. So the emission rate of particles can be expressed by the following equation:
@@
Eq. (186) indicates that our results conform to the unitary principle and support the conservation of information.
5 Quantum Correlations and Backreaction Effects
The tunneling formalism provides a semiclassical description of Hawking radiation, but a complete resolution of the information paradox requires incorporating quantum correlations and backreaction effects[51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. Here, we extend our analysis by integrating insights from recent advances in quantum gravity and non-thermal modifications.
5.1 Quantum correlations in sequential emissions
As emphasized in [51], the absence of correlations between successive emissions in the original Parikh-Wilczek framework limits its ability to address unitarity. To probe this, we calculate the joint tunneling probability for two particles with energies $E_{1}$ and $E_{2}$:
$\Gamma(E_{1},E_{2})\sim e^{\Delta S(M\to M-E_{1}-E_{2})}.$ (187)
contrasted with the uncorrelated case $\Gamma(E_{1})\Gamma(E_{2})\sim e^{\Delta S(M\to M-E_{1})}e^{\Delta S(M-E_{1}\to M-E_{1}-E_{2})}$. For Schwarzschild de Sitter black holes, we find:
$\Gamma(E_{1},E_{2})\neq\Gamma(E_{1})\Gamma(E_{2}),~{}~{}~{}~{}\text{if}~{}\frac{\partial^{2}S}{\partial M^{2}}\neq 0.$ (188)
indicating entropy-driven correlations. This aligns with [51]’s conclusion that self-gravity induces effective correlations, though not equivalent to microscopic entanglement. Such correlations partially resolve the "thermal spectrum paradox" but fall short of guaranteeing full unitarity — a gap likely bridged only by including subleading quantum gravity effects. While our semiclassical analysis confirms entropy conservation, full unitarity requires quantum correlations beyond thermodynamic entropy — a challenge for all tunneling-based approaches.
5.2 Backreaction and dynamic spacetime
The assumption of discrete mass transitions $M\to M-\omega$ n our original analysis approximates continuous evaporation. Following [52, 53], we refine this by modeling the black hole mass as a time-dependent function $M(t)=M_{0}-\int_{0}^{t}\Gamma(\omega)\omega\mathrm{d}\omega$, leading to a time-varying metric:
$f(r,t)=1-\frac{2M(t)}{r}+\frac{Q_{h}^{2}}{r^{2}}-\frac{\Lambda}{3}r^{2}.$ (189)
Numerical integration of the modified geodesic equations reveals that continuous mass loss suppresses late-time radiation $T_{h}\propto M(t)^{-1}$, consistent with Page curve behavior. This dynamic backreaction — omitted in static tunneling models — strengthens our conclusion that entropy reduction $\Delta S<0$ governs the terminal evaporation phase, mirroring unitary evolution. Dynamic backreaction and GUP effects predict testable deviations from Hawking’s thermal spectrum, notably in late-time evaporation rates $\frac{\mathrm{d}M}{\mathrm{d}t}\propto-M^{-2}$ and high-energy suppression.
5.3 Non-thermal modifications at high energy
For high-energy particles $\omega\sim M$, the WKB approximation breaks down. Adopting the generalized uncertainty principle (GUP) from [52], where $\Delta x\Delta p\geq\frac{\hbar}{2}(1+\beta(\Delta p)^{2})$, we derive a modified tunneling probability:
$\Gamma_{GUP}(\omega)\sim\exp\left(-\frac{\pi}{\hbar}\int_{r_{i}}^{r_{f}}\mathrm{d}r\frac{\sqrt{1+\beta(\partial_{r}S)^{2}}}{f(r)}\right).$ (190)
This introduces a UV cutoff $\beta>0$ suppressing high-energy emissions, akin to firewall scenarios. While our baseline results $\beta=0$ assume semiclassicality, this extension highlights how quantum gravity effects could imprint non-thermal signatures in future observations. These corrections align with dS/CFT proposals [40, 43, 42], where boundary unitarity might dynamically regulate bulk information flow.
6 Comparative Analysis with Alternative Approaches
The Parikh-Wilczek tunneling formalism provides a robust framework for studying semiclassical radiation processes in dynamic spacetime backgrounds. However, to contextualize our results within the broader landscape of black hole information research, we systematically compare our methodology with three prominent alternative approaches.
The Hamilton-Jacobi method calculates particle emission rates by solving the relativistic Hamilton-Jacobi equation for imaginary time trajectories [62, 63, 64]. While this approach shares the tunneling paradigm with our work, it inherently assumes static background metrics, neglecting self-gravitational effects — a critical limitation addressed by the Parikh-Wilczek formalism through energy-conserving backreacted geometries. Our analysis of horizon displacement $r_{+}\to r_{+}-\delta r$ under particle emission explicitly demonstrates how self-gravity modifies both the emission spectrum $\Gamma\sim e^{\Delta S}$ and entropy evolution, a feature absent in static-background approximations.
Recent advances in AdS/CFT derive unitary Page curves through island contributions to entanglement entropy [8, 9]. While these results resolve the paradox in AdS black holes, their extension to dS spacetime remains conjectural. Crucially, our entropy-driven formulation $\Delta S_{BH}=-2\mathrm{Im}S$ aligns thermodynamically with the Page curve’s late-time entropy decrease, yet differs fundamentally by relying on semiclassical geometry rather than replica wormholes. This highlights a key distinction: our approach captures coarse-grained thermodynamic irreversibility, while fine-grained unitarity requires additional quantum correlations — a gap bridged neither by tunneling nor entanglement entropy methods alone.
Generalized uncertainty principle (GUP) models predict modified dispersion relations that alter tunneling probabilities. For instance, GUP-induced $\Delta x\Delta p\geq\frac{\hbar}{2}(1+\beta(\Delta p)^{2})$ would perturb the WKB phase integral. Our formalism naturally accommodates such extensions through momentum-dependent metric corrections $g_{\mu\nu}\to g_{\mu\nu}(p)$, whereas traditional thermal spectrum analyses remain insensitive to Planck-scale effects. This adaptability positions the Parikh-Wilczek method as a versatile tool for probing quantum gravity imprints in de Sitter black hole radiation [65, 66, 67]. Non commutative geometry modifies the spatiotemporal structure by introducing coordinate non commutativity, thereby affecting the tunneling behavior of particles in quantum field theory[68, 69, 70, 56]. Non-commutative geometry introduces non thermal corrections to radiation spectra (such as energy dependent cutoff) by breaking the locality of classical spacetime, while preserving the macroscopic manifestation of information conservation. Non-commutative parameter $\theta$ also perturb the WKB phase integral $f(r)\to\tilde{f}(r,\theta)$.
In order to compare the differences among these methods more intuitively, we have summarized the characteristics of these methods in Table 1.
Table 1: Methodological Comparison
Aspect Parikh-Wilczek Hamilton-Jacobi
Background Dynamics Dynamic (self-gravity included) Static
Information Carrier Thermodynamic entropy ($\Delta S$) Tunneling phase
de Sitter Compatibility Explicit Limited to asymptotically flat
Microscopic Insights Coarse-grained None
Aspect AdS/CFT + Islands GUP-Based Models
Background Dynamics Euclidean path integrals Static with momentum corrections
Information Carrier Entanglement entropy Modified dispersion relations
de Sitter Compatibility Conjectural (dS/CFT required) Requires dS extension
Microscopic Insights Fine-grained (replica wormholes) Planck-scale modifications
We summarize our comparison of these methods as follows. The Parikh-Wilczek method uniquely balances dynamic spacetime treatment (essential for de Sitter cosmology) and analytical tractability, making it ideal for probing multi-parameter black holes (charged, rotating, regular). While entanglement entropy frameworks [8, 9] provide deeper microscopic insights, our results offer complementary thermodynamic consistency checks for de Sitter holography. Future synergies could emerge by hybridizing tunneling calculations with GUP corrections to quantify quantum gravity’s role in information recovery.